OpenStudy (anonymous):

Can someone answer this and explain it. noah and kate are shopping for a new guitar strings in a music store and noah buys 2 packs of strings. kate buys 2 packs of strings and a music book. the book cost $16. their total cost is $72. how much is one pack of strings? solve using multi-step equation please

OpenStudy (anonymous):

let b represent books and s represent strings 2s (from noah) + 2s (from kate)+b=72 if b=16 then 4s=72-16 s=14

OpenStudy (anonymous):

thank you!!

OpenStudy (aravindg):

\[\huge \text{WELCOME TO OPENSTUDY !}\]